# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### v2

@@ E.2 Two moons: targeting the posterior directly vs targeting the likelihood
In the main paper we discussed the two possibilites of targeting the posterior $\tilde{p}_{r+1}(\bm{\theta})=q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ by first obtaining an estimate $q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})$ of the likelihood and then sampling via MCMC from $\tilde{p}_{r+1}(\bm{\theta})$, as opposed to the option of using the GLLiM surrogate posterior $q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$ to update the posterior $\tilde{p}_{r+1}(\bm{\theta})=\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$, and then sampling from the latter via MCMC. We hinted at the fact that it is preferable to follow the first route, which we used to obtain all our results, and here we provide results in support of our recommendation. Figure 7 shows the C2ST metric and the Wasserstein distance as a function of the number of model simulations for different SeMPLE settings, including different choices for the covariance matrices $\tilde{\bm{\Sigma}}_{k}$ (“isotropic” or “full unconstrained”) of the mixture models. We first considered the “isotropic” covariances case (specified in xLLiM via the option cstr\$Sigma="i"), which means that all the $K$ covariance matrices are proportional to the diagonal matrix, however these matrices are all different between the $K$ components. With this option, we ran SeMPLE by targeting $\tilde{p}_{r+1}(\bm{\theta})=q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ (“likelihood; isotropic” in the legend of Figure 7) and also ran SeMPLE by targeting $\tilde{p}_{r+1}(\bm{\theta})=\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$ (“posterior; isotropic” in the legend of Figure 7). Targeting $\tilde{p}_{r+1}(\bm{\theta})=q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ is clearly the preferred route according to C2ST. This is less clear by looking at the Wasserstein distance, however the computation of this particular metric may not be accurate with multimodal targets (while we prefer the latter for unimodal targets), and we refer the reader to section 5.5 in the main paper for a discussion about metrics to evaluate posterior inference. Additionally, we explored the results of assuming full matrices $\tilde{\bm{\Sigma}}_{k}$ without constraints (specified in xLLiM via the option cstr\$Sigma=" "). This setting is run together with our (recommended) method of sampling from $\tilde{p}_{r+1}(\bm{\theta})=q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ (“likelihood; unconstrained” in the legend of Figure 7). This SeMPLE configuration produces the best results. This may not be surprising, as for two-moons $\bm{y}_{0}$ has dimension 2, therefore the covariances $\tilde{\bm{\Sigma}}_{k}$ have dimensions $2\times 2$, thus not inducing a large number of hyperparameters to estimate via EM. Notice that the C2ST value, when using our recommended method but with isotropic covariance matrices, is still lower than C2ST as obtained with SNPE-C (see the main paper). Regarding the performance of the MCMC algorithm, Figure 6 shows that the acceptance rate is noticeably higher with the “likelihood” option rather than the “posterior” one, which strengthens our choice. While a higher acceptance rate does not denote, per-se, any special ability for an algorithm to efficiently explore the posterior, however this fact coupled with the already displayed high quality inference provided by SeMPLE for this example, means that Figure 6 further shows that we can obtain a larger number of representative posterior samples when targeting $\tilde{p}_{r+1}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$.
@@ E.3 Two moons: selection of $K$ via BIC
In the main paper, we have considered the inference results based on a total of $10^{4}$ model simulations, uniformly distributed across $R=10$ rounds for SNPE-C and SNL (as per SBIBM defaults), and $R=4$ rounds for SeMPLE. Results showed that SeMPLE performed better and in particular that SNL (which is also an MCMC-based procedure, similarly to SeMPLE) was returning the worst performance. Here we show the performance of both SNPE-C and SNL with $R=4$, again using a total of $10^{4}$ model simulations for each method. That is, for each method we show the inference obtained after 2500, 5000, 7500 and 10000 model simulations. Figures 9-18 shows scatter plots (colored by density value) of the posterior samples from the last round (r=4) of each algorithm and for each of the 10 observed data sets in SBIBM. SeMPLE consistently matches the reference posterior, unlike SNL and SNPE-C.
@@
Figure 9: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 1  in SBIBM.
@@
Figure 10: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 2  in SBIBM.
@@
Figure 11: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 3  in SBIBM.
@@
Figure 12: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 4  in SBIBM.
@@
Figure 13: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 5  in SBIBM.
@@
Figure 14: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 6  in SBIBM.
@@
Figure 15: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 7  in SBIBM.
@@
Figure 16: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 8  in SBIBM.
@@
Figure 17: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 9  in SBIBM.
@@
Figure 18: Two Moons. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 10  in SBIBM.
@@
Moreover, it is even more interesting to note that SeMPLE performs already well after a single round, see Figures 19-28 showing posterior samples from every round of each algorithm and for each of the 10 observed data sets. While SNL improves across rounds, this is not the case for SNPE-C.
@@
Figure 19: Two Moons, inference using observed dataset number 1  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 20: Two Moons, inference using observed dataset number 2  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 21: Two Moons, inference using observed dataset number 3  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 22: Two Moons, inference using observed dataset number 4  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 23: Two Moons, inference using observed dataset number 5  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 24: Two Moons, inference using observed dataset number 6  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 25: Two Moons, inference using observed dataset number 7  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 26: Two Moons, inference using observed dataset number 8  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 27: Two Moons, inference using observed dataset number 9  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 28: Two Moons, inference using observed dataset number 10  from SBIBM. Posterior samples from each algorithm round, starting from the first (upper row) to the fourth (bottom row) round.
@@
Figure 29 shows C2ST and runtime values in the case when $R=4$ rounds are used for all algorithms. Compared to the results in the main paper with $R=10$ rounds, SNPE-C and SNL perform slightly worse with $R=4$ rounds in terms of the C2ST metric. The SNPE-C and SNL runtimes are however reduced, with SNPE-C having essentially identical runtime as SeMPLE, although with significantly higher variance. SNL is approximately twice as slow as the other algorithms.
@@
Figure 29: Two Moons. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.
@@ F.1 Multiple hyperboloid model definition
Since the multiple hyperboloid model is not implemented in SBIBM, there exists no observed data sets or reference posterior samples in SBIBM. To produce an observed data set $\bm{y}_{0}$, the model simulator $p(\bm{y}\,|\,\bm{\theta}^{*})$ was run with the true parameter value $\bm{\theta}^{*}=(1.5,1)$. Since the posterior shape varies widely depending on the true parameter value, and to keep the connection to the experiment in Forbes et al. (2022), only a single such observed data set was produced. Instead, the metric results were averaged over 10 independent runs with this same observed data set. To produce a reference posterior sample corresponding to the produced observed data set, the Metropolis-Hastings algorithm was used. The sample size of the reference posterior sample was set to $10^{4}$.
@@ F.3 Multiple hyperboloid model: results with $R=4$
In the main paper we have considered the inference results based on a total of $4\times 10^{4}$ model simulations, uniformly distributed across $R=10$ rounds for SNPE-C and SNL (as per SBIBM defaults), and $R=4$ rounds for SeMPLE. Results showed that SeMPLE performed better and in particular that SNL (which is also an MCMC-based procedure, similarly to SeMPLE) was returning the worst performance. Here we show the performance of both SNPE-C and SNL with $R=4$, again using a total of $4\times 10^{4}$ model simulations for each method. Figures 31-33 show posterior samples from the last round (r=4) of each algorithm for each of the 10 independent algorithm runs. We can clearly notice that SNL fails completely and therefore, for ease of reading, the SNL marginal posteriors are not reported. However the differences between SNPE-C and SeMPLE are more subtle. It appears that occasionally SeMPLE samples from central areas, where there should not be any posterior mass, while SNPE-C seems to undersample some of the branches. Therefore, we also plot the corresponding kernel-smoothed marginals in Figures 31-33. There, it appears that the marginals from SeMPLE are overall closer to the reference marginals, which is in agreement with the better C2ST attained by SeMPLE (see the main text).
Figure 34 shows C2ST and runtime values in the case when $R=4$ rounds for all algorithms. The C2ST results of SNPE-C and SNL are very similar with $R=4$, meaning that SNL fails completely and SeMPLE still performs much better than SNPE-C. In terms of runtimes SNPE-C is still about 7 times slower than SeMPLE, disregarding SNL as the inference from this method is not useful.
@@
Figure 34: Hyperboloid. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.

### v3

@@ E.2 Two moons: targeting the posterior directly vs targeting the likelihood
In the main paper we discussed the two possibilities of targeting the posterior $\tilde{p}_{r+1}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ by first obtaining an estimate $q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})$ of the likelihood and then sampling via MCMC from $\tilde{p}_{r+1}(\bm{\theta})$, as opposed to the option of using the GLLiM surrogate posterior $q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$ to update the posterior $\tilde{p}_{r+1}(\bm{\theta})=\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$, and then sampling from the latter via MCMC. We hinted at the fact that it is preferable to follow the first route, which we used to obtain all our results, and here we provide results in support of our recommendation. Figure 8 shows the C2ST metric and the Wasserstein distance as a function of the number of model simulations for different SeMPLE settings, including different choices for the covariance matrices $\tilde{\bm{\Sigma}}_{k}$ (“isotropic” or “full unconstrained”) of the mixture models. We first considered the “isotropic” covariances case (specified in xLLiM via the option cstr\$Sigma="i"), which means that all the $K$ covariance matrices are proportional to the diagonal matrix, however these matrices are all different between the $K$ components. With this option, we ran SeMPLE by targeting $\tilde{p}_{r+1}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ (“likelihood; isotropic” in the legend of Figure 8) and also ran SeMPLE by targeting $\tilde{p}_{r+1}(\bm{\theta})=\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$ (“posterior; isotropic” in the legend of Figure 8). Targeting $\tilde{p}_{r+1}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ is clearly the preferred route according to C2ST. This is less clear by looking at the Wasserstein distance, however the computation of this particular metric may not be accurate with multimodal targets (while we recommend to use Wasserstein distances for unimodal targets), and we refer the reader to section 5.6 in the main paper for a discussion about metrics to evaluate posterior inference. Additionally, we explored the results of assuming full matrices $\tilde{\bm{\Sigma}}_{k}$ without constraints (specified in xLLiM via the option cstr\$Sigma=" "). This setting is run together with our (recommended) method of sampling from $\tilde{p}_{r+1}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ (“likelihood; unconstrained” in the legend of Figure 8). This SeMPLE configuration produces the best results. This may not be surprising, as for two-moons $\bm{y}_{0}$ has dimension 2, therefore the covariances $\tilde{\bm{\Sigma}}_{k}$ have dimensions $2\times 2$, thus not inducing a large number of parameters $\tilde{\bm{\phi}}$ to estimate via EM. Notice that the C2ST value, when using our recommended method but with isotropic covariance matrices, is still lower than C2ST as obtained with SNPE-C (see the main paper). Regarding the performance of the MCMC algorithm, Figure 7 shows that the acceptance rate is noticeably higher with the “likelihood” option rather than the “posterior” one, which strengthens our choice. While a higher acceptance rate does not denote, per-se, any special ability for an algorithm to efficiently explore the posterior, however this fact coupled with the already displayed high quality inference provided by SeMPLE for this example, means that Figure 7 further shows that we can obtain a larger number of representative posterior samples when targeting $\tilde{p}_{r+1}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$.
@@ E.3 Two moons: selection of $K$ via BIC
In the main paper, we have considered the inference results based on a total of $10^{4}$ model simulations, uniformly distributed across $R=10$ rounds for SNPE-C and SNL (as per SBIBM defaults), and $R=4$ rounds for SeMPLE. Results showed that SeMPLE performed better and in particular that SNL (which is also an MCMC-based procedure, similarly to SeMPLE) was returning the worst performance. Here we show the performance of both SNPE-C and SNL with $R=4$, again using a total of $10^{4}$ model simulations for each method. That is, for each method we show the inference obtained after 2500, 5000, 7500 and 10000 model simulations. Figures 10-19 shows scatter plots (colored by density value) of the posterior samples from the last round (r=4) of each algorithm and for each of the 10 observed data sets in SBIBM. SeMPLE consistently matches the reference posterior, unlike SNL and SNPE-C.
@@
Figure 10: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 1  in SBIBM.
@@
Figure 11: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 2  in SBIBM.
@@
Figure 12: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 3  in SBIBM.
@@
Figure 13: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 4  in SBIBM.
@@
Figure 14: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 5  in SBIBM.
@@
Figure 15: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 6  in SBIBM.
@@
Figure 16: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 7  in SBIBM.
@@
Figure 17: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 8  in SBIBM.
@@
Figure 18: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 9  in SBIBM.
@@
Figure 19: Two Moons using $R=4$ rounds for every method. Density scatter plot of posterior samples from the last round (r=4) of each algorithm. Observed data set number 10  in SBIBM.
@@
Moreover, it is even more interesting to note that SeMPLE performs already well after a single round, see Figures 20-29 showing posterior samples from every round of each algorithm and for each of the 10 observed data sets. While SNL improves across rounds, this is not the case for SNPE-C.
@@
Figure 20: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 1  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 21: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 2  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 22: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 3  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 23: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 4  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 24: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 5  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 25: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 6  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 26: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 7  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 27: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 8  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 28: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 9  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 29: Two Moons using $R=4$ rounds for every method: inference using observed dataset number 10  from SBIBM. Posterior samples from each algorithm round, starting from the first round (upper row) to the fourth (bottom row).
@@
Figure 30 shows C2ST and runtime values in the case when $R=4$ is used for all algorithms. Compared to the results in the main paper with $R=10$ rounds, SNPE-C and SNL perform slightly worse with $R=4$ rounds in terms of the C2ST metric. The SNPE-C and SNL runtimes are however reduced, with SNPE-C having essentially identical runtime as SeMPLE, although with significantly higher variance. SNL is approximately twice as slow as the other algorithms.
@@
Figure 30: Two Moons using $R=4$ rounds for every method. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.
@@ F.1 Multiple hyperboloid model definition
Since the multiple hyperboloid model is not implemented in SBIBM, there exists no observed data sets or reference posterior samples in SBIBM. To produce an observed data set $\bm{y}_{0}$, the model simulator $p(\bm{y}\,|\,\bm{\theta}^{*})$ was run with the true parameter value $\bm{\theta}^{*}=(1.5,1)$. Since the posterior shape varies widely depending on the true parameter value, and to keep the connection to the experiment in Forbes et al. (2022), only a single observed data set was produced. Instead, the metric results were averaged over 10 independent runs with this same observed data set. To produce a reference posterior sample corresponding to the produced observed data set, the Metropolis-Hastings algorithm was used. The sample size of the reference posterior sample was set to $10^{4}$.
@@ F.3 Multiple hyperboloid model: results with $R=4$
In the main paper we have considered the inference results based on a total of $4\times 10^{4}$ model simulations, uniformly distributed across $R=10$ rounds for SNPE-C and SNL (as per SBIBM defaults), and $R=4$ rounds for SeMPLE. Results showed that SeMPLE performed better and in particular that SNL (which is also an MCMC-based procedure, similarly to SeMPLE) was returning the worst performance. Here we show the performance of both SNPE-C and SNL with $R=4$, again using a total of $4\times 10^{4}$ model simulations for each method. Figures 32-34 show posterior samples from the last round (r=4) of each algorithm for each of the 10 independent algorithm runs. We can clearly notice that SNL fails completely and therefore, for ease of reading, the SNL marginal posteriors are not reported. However, the differences between SNPE-C and SeMPLE are more subtle. It appears that occasionally SeMPLE samples from central areas, where there should not be any posterior mass, while SNPE-C seems to undersample some of the branches. Therefore, we also plot the corresponding kernel-smoothed marginals in Figures 32-34. There, it appears that the marginals from SeMPLE are overall closer to the reference marginals, which is in agreement with the better C2ST attained by SeMPLE (see the main text).
Figure 35 shows C2ST and runtime values in the case when $R=4$ rounds for all algorithms. The C2ST results of SNPE-C and SNL are very similar with $R=4$, meaning that SNL fails completely and SeMPLE still performs much better than SNPE-C. In terms of runtime, SNPE-C is still about 7 times slower than SeMPLE, disregarding SNL as the inference from this method is not useful.
@@
Figure 35: Hyperboloid. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with the same data vs the number of model simulations. Shaded bands enclose the min and max values.